40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5657     Entry time: Wed Oct 12 18:54:02 2011
Author: Katrin 
Type: Update 
Category: Green Locking 
Subject: 60 Hz oscillation due to broken BNC cable 

There was a 60 Hz and 120 Hz oscillation on the green PDH photo diode output. After a long search, I could identify that

the source was a broken BNC cable which was connected to the photo diode. I exchanged that BNC cable and the 60 Hz

and 120 Hz are gone :-)

With the new cable the PD output was less noisy so that it was easier to achieve a better alignment of the light to the cavity.

The reflected power could be reduced from 40% to 30%. For perfect alignment the reflected power would be 20%.

ELOG V3.1.3-